OpenStudy (anonymous):

Evaluate 8(x - 3y) when x = 2 and y = 4 a. 112 b. 80 c. -80 d. -16

OpenStudy (anonymous):

blah;(

Directrix (directrix):

@khrisangel What is the missing symbol in 8(x 3y)? (between x and 3y)

OpenStudy (anonymous):

8(x-3y)

OpenStudy (aripotta):

i would distribute first, then substitute

OpenStudy (aripotta):

or not, it's easy both ways

Directrix (directrix):

8(x - 3y) = 8[ (2) - 3(4) ] = 8 [ 2 - 12 ] = 8 [ -10 ] = - 80